Really need help thanks, 2 decimals

A stock just paid a dividend of $1.82. The dividend is expected to grow at 27.55% for five years and then grow at 3.60% thereafter. The required return on the stock is 12.23%. What is the value of the stock?
Answer #1

D1 = 1.82(1.2755) = $2.32

D2 = 1.82(1.2755)2 = $2.96

D3 = 1.82(1.2755)3 = $3.78 + ()

D4 = 1.82(1.2755)4 = $4.82

D5 = 1.82(1.2755)5 = $6.14

Horizon Value = 6.14(1.036)/(0.1223 - 0.036)

Horizon Value = $73.71

Using DDM Model,

Stock Price = 2.32/(1.223) + 2.96/(1.1223)2 + 3.78/(1.1223)3 + 4.82/(1.1223)4 + (73.71 + 6.14)/(1.1223)5

Stock Price = $54.98